# When UAV Meets Federated Learning: Latency Minimization via Joint Trajectory Design and Resource Allocation

### v1

@@ IV-E Communication and Computing Resource Allocation
Please refer to Appendix B. ∎
@@ IV-F Algorithm Analysis
The proposed algorithm for solving problem (24) is provided in Algorithm 1, which alternately optimizes the RB allocation, UAV trajectory, and communication and computation resource allocation in an iterative manner until the objective value converges or the maximum iteration number is reached. The following two propositions analyze the convergence and complexity of Algorithm 1.
Proposition 2.
Proof.
Please refer to Appendix C. ∎
@@
Please refer to Appendix D. ∎
@@ V Numerical Results
TABLE I: Main parameter settings.
@@ V-C Overall Latency
Fig. 4 shows the total latency of all IoT devices versus the number of RBs. We can observe that compared to other benchmark schemes, the proposed scheme achieves the lowest latency. The reason is that the proposed scheme can simultaneously optimize the allocation of RBs, transmit power, computing frequency, and UAV trajectory, based on the communication quality between the UAV and each IoT device. As the number of RBs increases, the latency for all schemes decreases, which can be attributed to the increased bandwidth resources for improving the data rates among the UAV and IoT devices.
@@ VI Conclusion
Appendix A Proof of the Theorem 2
@@
Appendix B Proof of the Proposition 1
@@
Appendix C Proof of the Proposition 2
@@
Appendix D Proof of the Proposition 3
@@
The computational complexity of Algorithm 1 is primarily composed of solving problems (34), (50), and (66). In problem (34), the complexity for optimizing the variables $\{\boldsymbol{\alpha},\boldsymbol{\eta}\}$ is $(MUN+N)^{3.5}$, where $MUN+N$ is the number of variables. In problem (50), the complexity for optimizing the variables $\{\mathbf{Q},\boldsymbol{\eta},\mathbf{d},\mathbf{R}\}$ is $(3MN+N)^{3.5}$, where $3MN+N$ is the number of variables. In problem (66), the complexity for optimizing the variables $\{\mathbf{f},\mathbf{p},\boldsymbol{\eta},\boldsymbol{\gamma},\mathbf{\Xi}\}$ is $(7MN+2N)^{3.5}$, where $MUN+N$ is the number of variables. Therefore, the total computational complexity is $\Omega=\mathcal{O}(\mathcal{L}((MUN+N)^{3.5}+(3MN+3N)^{3.5})\quad+(7MN+2N)^{3.5})\log(1/\epsilon_{a}))$. ∎

### v2

@@ IV-E Communication and Computing Resource Allocation
Please refer to Appendix D. ∎
@@ IV-F Algorithm Analysis
The proposed algorithm for solving problem (24) is provided in Algorithm 1, which alternately optimizes the RB allocation, the UAV trajectory, and communication and computation resource allocation in an iterative manner until the objective value converges or the maximum iteration number is reached. The following two propositions analyze the convergence and complexity of Algorithm 1.
@@
Please refer to Appendix F. ∎
@@ V Numerical Results
TABLE II: Main parameter settings.
@@
The density of the Poisson distribution, $\rho_{\mathrm{sensed}}$ and $\rho_{\mathrm{dropped}}$ $50$
The data size of local model updates, $d_{m}[n]$ $[3.5-4.5]*10^{5}\rm{bits}$
@@ V-C Overall Latency
Fig. 5 shows the total latency of all IoT devices versus the number of RBs. We can observe that compared to other benchmark schemes, the proposed scheme achieves the lowest latency. The reason is that the proposed scheme can simultaneously optimize the allocation of RBs, transmit power, computing frequency, and the UAV trajectory, based on the communication quality between the UAV and each IoT device. As the number of RBs increases, the latency for all schemes decreases, which can be attributed to the increased bandwidth resources for improving the data rates among the UAV and IoT devices.
@@ VI Conclusion
Appendix A Proof of the Proposition 1
Proof.
During local training, the local optimization problem of the $m$-th IoT device satisfies,
~$\tilde{\mathcal{L}}(\boldsymbol{\theta}_{m}^{0};\mathcal{D}_{m})\geq\tilde{\mathcal{L}}(\boldsymbol{\theta}_{m}^{1};\mathcal{D}_{m})\geq\cdots\geq\tilde{\mathcal{L}}(\boldsymbol{\theta}_{m}^{r};\mathcal{D}_{m}),$ (68)
where the time slot index $n$ are omitted for simplicity. To guide convergence, the iteration of local training with accuracy $\epsilon_{\mathrm{Local}}$ should satisfy
~$\begin{split}\tilde{\mathcal{L}}(\boldsymbol{\theta}_{m}^{r};\mathcal{D}_{m})&-\tilde{\mathcal{L}}(\boldsymbol{\theta}_{m}^{*};\mathcal{D}_{m})\\ &\leq\epsilon_{\mathrm{Local}}\left(\tilde{\mathcal{L}}(\boldsymbol{\theta}_{m}^{0};\mathcal{D}_{m})-\tilde{\mathcal{L}}(\boldsymbol{\theta}_{m}^{*};\mathcal{D}_{m})\right),\end{split}$ (69)
where $\boldsymbol{\theta}_{m}^{*}$ denotes the optimal parameter of the $m$-th IoT device. From (5), it follows that [32]
~$\begin{split}\tilde{\mathcal{L}}(\boldsymbol{\theta}_{m}^{r};\mathcal{D}_{m})\leq\tilde{\mathcal{L}}(\boldsymbol{\theta}_{m}^{r-1};\mathcal{D}_{m})-\upsilon\bar{\mathcal{L}}_{m}^{r-1}+\frac{L\upsilon^{2}}{2}\bar{\mathcal{L}}_{m}^{r-1},\end{split}$ (70)
where ¯$\bar{\mathcal{L}}_{m}^{r-1}=\|\nabla\tilde{\mathcal{L}}(\boldsymbol{\theta}_{m}^{r-1};\mathcal{D}_{m})\|^{2}$. According to Lemma 5 in [32], we have
$\Upsilon\left(\tilde{\mathcal{L}}(\boldsymbol{\theta}_{m}^{r};\mathcal{D}_{m})-\tilde{\mathcal{L}}(\boldsymbol{\theta}_{m}^{*};\mathcal{D}_{m})\right)\leq\bar{\mathcal{L}}_{m}^{r}.$ (71)
Under Assumptions 1 and 2, by combining (68), (69), (70), and (71), we obtain
~$\begin{split}\tilde{\mathcal{L}}(&\boldsymbol{\theta}_{m}^{r};\mathcal{D}_{m})-\tilde{\mathcal{L}}(\boldsymbol{\theta}_{m}^{*};\mathcal{D}_{m})\leq\tilde{\mathcal{L}}(\boldsymbol{\theta}_{m}^{r-1};\mathcal{D}_{m})-\tilde{\mathcal{L}}(\boldsymbol{\theta}_{m}^{*};\mathcal{D}_{m})\\ &\quad-\left(\upsilon+\frac{L\upsilon^{2}}{2}\right)\Upsilon\left(\tilde{\mathcal{L}}(\boldsymbol{\theta}_{m}^{r-1};\mathcal{D}_{m})-\tilde{\mathcal{L}}(\boldsymbol{\theta}_{m}^{*};\mathcal{D}_{m})\right)\\ &\leq\left(1-\frac{\upsilon\Upsilon(2-L\upsilon)}{2}\right)\left(\tilde{\mathcal{L}}(\boldsymbol{\theta}_{m}^{r-1};\mathcal{D}_{m})-\tilde{\mathcal{L}}(\boldsymbol{\theta}_{m}^{*};\mathcal{D}_{m})\right)\\ &\leq\left(1-\frac{\upsilon\Upsilon(2-L\upsilon)}{2}\right)^{r}\left(\tilde{\mathcal{L}}(\boldsymbol{\theta}_{m}^{0};\mathcal{D}_{m})-\tilde{\mathcal{L}}(\boldsymbol{\theta}_{m}^{*};\mathcal{D}_{m})\right)\\ &\leq\mathrm{e}^{\left(-\frac{\upsilon\Upsilon(2-L\upsilon)}{2}\right)r}\left(\tilde{\mathcal{L}}(\boldsymbol{\theta}_{m}^{0};\mathcal{D}_{m})-\tilde{\mathcal{L}}(\boldsymbol{\theta}_{m}^{*};\mathcal{D}_{m})\right).\end{split}$ (72)
To ensure the inequality (69), the lower bound of the local training rounds $\check{R}$ should satisfy the following condition, i.e.,
$\begin{split}\mathrm{e}^{\left(-\frac{\upsilon\Upsilon(2-L\upsilon)}{2}\right)\check{R}}=\epsilon_{\mathrm{Local}}.\end{split}$ (73)
By performing a simple transformation to the inequality (73), we can obtain (28). ∎
Appendix B Proof of the Proposition 2
Proof.
Similarly, the iteration round of the UAV model aggregation should satisfy
$\begin{split}{\mathcal{L}}(\boldsymbol{\theta}[n])-{\mathcal{L}}(\boldsymbol{\theta}^{*})\leq\epsilon_{\mathrm{UAV}}\left({\mathcal{L}}(\boldsymbol{\theta}[0])-{\mathcal{L}}(\boldsymbol{\theta}^{*})\right).\end{split}$ (74)
From (4), (7), and Assumptions 1 and 2, it follows that [32]
$\begin{split}{\mathcal{L}}(&\boldsymbol{\theta}[n])\leq{\mathcal{L}}(\boldsymbol{\theta}[n-1])+\frac{1}{M}\sum_{m=1}^{M}\nabla{\mathcal{L}}(\boldsymbol{\theta}[n-1])^{\mathsf{T}}\tilde{\theta}_{m}[n-1]\\ &\quad\quad\quad\ +\frac{L}{2M^{2}}\left\|\sum_{m=1}^{M}\tilde{\theta}_{m}[n-1]\right\|^{2}\\ &\leq{\mathcal{L}}(\boldsymbol{\theta}[n-1])+\frac{1}{M\varrho}\sum_{m=1}^{M}\Bigg{(}\tilde{\mathcal{L}}(\boldsymbol{\theta}_{m}[n-1];D_{m}[n-1])\\ &\quad-{\mathcal{L}}(\boldsymbol{\theta}_{m}[n-1];D_{m}[n-1])-\frac{\Upsilon}{2}\left\|\tilde{\theta}_{m}[n-1]\right\|^{2}\Bigg{)}\\ &\quad+\frac{L}{2M^{2}}\left\|\sum_{m=1}^{M}\tilde{\theta}_{m}[n-1]\right\|^{2}\\ &\leq{\mathcal{L}}(\boldsymbol{\theta}[n-1])+\frac{1}{M\varrho}\sum_{m=1}^{M}\Bigg{(}\tilde{\mathcal{L}}(\boldsymbol{\theta}_{m}[n-1];D_{m}[n-1])\\ &\quad-{\mathcal{L}}(\boldsymbol{\theta}_{m}[n-1];D_{m}[n-1])-\frac{\Upsilon-L\varrho}{2}\left\|\tilde{\theta}_{m}[n-1]\right\|^{2}\Bigg{)}\\ &\leq{\mathcal{L}}(\boldsymbol{\theta}[n-1])-\frac{1}{M\varrho}\sum_{m=1}^{M}\Bigg{(}\frac{\Upsilon-L\varrho}{2}\left\|\tilde{\theta}_{m}[n-1]\right\|^{2}\\ &\quad+\frac{\Upsilon(1-\epsilon_{\mathrm{local}})}{2}\left\|\tilde{\theta}_{m}^{*}[n-1]\right\|^{2}\Bigg{)}.\end{split}$ (75)
From Assumptions 1 and 2, we have
$\frac{1}{L^{2}}\left\|\nabla{\mathcal{L}}(\boldsymbol{\theta}_{m}^{*}[n];D_{m}[n])-\nabla{\mathcal{L}}(\boldsymbol{\theta}[n];D_{m}[n])\right\|^{2}\leq\left\|\tilde{\theta}_{m}^{*}[n]\right\|^{2}.$ (76)
Combining (75) and (76), we have
$\begin{split}{\mathcal{L}}(\boldsymbol{\theta}[n])&\leq{\mathcal{L}}(\boldsymbol{\theta}[n-1])-\frac{1}{M\varrho}\sum_{m=1}^{M}\Bigg{(}\frac{\Upsilon-L\varrho}{2}\left\|\tilde{\theta}_{m}[n-1]\right\|^{2}\\ &\quad+\frac{\Upsilon(1-\epsilon_{\mathrm{local}})}{2L^{2}}\left\|\nabla{\mathcal{L}}(\boldsymbol{\theta}_{m}^{*}[n-1];D_{m}[n-1])\right.\\ &\quad\left.-\nabla{\mathcal{L}}(\boldsymbol{\theta}[n-1];D_{m}[n-1])\right\|^{2}\Bigg{)}.\end{split}$ (77)
When the local training problem (4) reach the optimal solution, we have the following condition,
$\nabla{\mathcal{L}}(\boldsymbol{\theta}_{m}^{*}[n];D_{m}[n])-\nabla{\mathcal{L}}(\boldsymbol{\theta}[n];D_{m}[n])=-\varrho\nabla{\mathcal{L}}(\boldsymbol{\theta}[n]).$ (78)
Substituting (78) into (77), we have
$\begin{split}{\mathcal{L}}(\boldsymbol{\theta}[n])&\leq{\mathcal{L}}(\boldsymbol{\theta}[n-1])-\frac{1}{M\varrho}\sum_{m=1}^{M}\Bigg{(}\frac{\Upsilon-L\varrho}{2}\left\|\tilde{\theta}_{m}[n-1]\right\|^{2}\\ &\quad+\frac{\Upsilon\varrho(1-\epsilon_{\mathrm{local}})}{2L^{2}}\left\|\nabla{\mathcal{L}}(\boldsymbol{\theta}[n-1])\right\|^{2}\Bigg{)}.\\ &\leq{\mathcal{L}}(\boldsymbol{\theta}[n-1])-\frac{1}{M\varrho}\sum_{m=1}^{M}\Bigg{(}\frac{\Upsilon-L\varrho}{2}\left\|\tilde{\theta}_{m}[n-1]\right\|^{2}\\ &\quad+\frac{\Upsilon^{2}\varrho(1-\epsilon_{\mathrm{local}})}{2L^{2}}\left({\mathcal{L}}(\boldsymbol{\theta}[n-1])-\mathcal{L}(\boldsymbol{\theta}^{*})\right)\Bigg{)}.\end{split}$ (79)
Since we set $0\leq\varrho\leq\frac{\Upsilon}{L}$, i.e., $0\leq\Upsilon-L\varrho$, by subtracting $\mathcal{L}(\boldsymbol{\theta}^{*})$ from both sides of the inequality (79), (79) can be rewritten as
$\begin{split}{\mathcal{L}}&(\boldsymbol{\theta}[n])-\mathcal{L}(\boldsymbol{\theta}^{*})\leq{\mathcal{L}}(\boldsymbol{\theta}[n-1])-\mathcal{L}(\boldsymbol{\theta}^{*})\\ &\quad-\left(\frac{\Upsilon^{2}\varrho(1-\epsilon_{\mathrm{local}})}{2L^{2}}\right)({\mathcal{L}}(\boldsymbol{\theta}[n-1])-\mathcal{L}(\boldsymbol{\theta}^{*}))\\ &=\left(1-\frac{\Upsilon^{2}\varrho(1-\epsilon_{\mathrm{local}})}{2L^{2}}\right)({\mathcal{L}}(\boldsymbol{\theta}[n-1])-\mathcal{L}(\boldsymbol{\theta}^{*}))\\ &\leq\left(1-\frac{\Upsilon^{2}\varrho(1-\epsilon_{\mathrm{local}})}{2L^{2}}\right)^{n}({\mathcal{L}}(\boldsymbol{\theta}[0])-\mathcal{L}(\boldsymbol{\theta}^{*}))\\ &\leq\mathrm{e}^{-\frac{\Upsilon^{2}\varrho(1-\epsilon_{\mathrm{local}})}{2L^{2}}n}({\mathcal{L}}(\boldsymbol{\theta}[0])-\mathcal{L}(\boldsymbol{\theta}^{*})).\end{split}$ (80)
Similar to Appendix A, the following condition should hold for the lower bound of the global aggregation rounds $\check{N}$, i.e.,
$\mathrm{e}^{-\frac{\Upsilon^{2}\varrho(1-\epsilon_{\mathrm{local}})}{2L^{2}}\check{N}}=\epsilon_{\mathrm{UAV}}.$ (81)
Then, we can derive (29). ∎
Appendix C Proof of the Theorem 1
@@
Appendix D Proof of the Proposition 3
@@
Appendix E Proof of the Proposition 4
@@
Appendix F Proof of the Proposition 5
@@
The computational complexity of Algorithm 1 is primarily composed of solving problems (34), (50), and (66). In problem (34), the complexity for optimizing the variables $\{\boldsymbol{\alpha},\boldsymbol{\eta}\}$ is $(MUN+N)^{3.5}$, where $MUN+N$ is the number of variables. In problem (50), the complexity for optimizing the variables $\{\mathbf{Q},\boldsymbol{\eta},\mathbf{d},\mathbf{R}\}$ is $(3MN+N)^{3.5}$, where $3MN+N$ is the number of variables. In problem (66), the complexity for optimizing the variables $\{\mathbf{f},\mathbf{p},\boldsymbol{\eta},\boldsymbol{\gamma},\mathbf{\Xi}\}$ is $(7MN+2N)^{3.5}$, where $MUN+N$ is the number of variables. Therefore, the total computational complexity is $\Omega=\mathcal{O}(\mathcal{L}((MUN+N)^{3.5}+(3MN+3N)^{3.5})+(7MN+2N)^{3.5})\log(1/\epsilon_{a}))$. ∎